what is 5 2/3 divided by 4 1/2 =
Answer by Alan3354(69443)

what is 5 2/3 divided by 4 1/2 =
-----------
Change to improper fractions.
17/3 / 9/2
Invert the divisor and multiply
= 17/3 * 2/9
= 34/27